Quartz, SiO2, is the most abundant mineral found at Earth's surface and its unique properties make it one of the most useful natural substances. Quartz is a chemical compound consisting of one part silicon and two parts oxygen, what is its percentage composition?​

1
Expert's answer
2022-01-14T03:12:04-0500

Solution:

Quartz = SiO2

Relative atomic mass (Ar) of silicon (Si) = 28.086

Ar(O) = 15.999

Molar mass of SiO2 = Ar(Si) + 2 × Ar(O) = 28.086 + 2 × 15.999 = 60.084

Hence,

%Si = [Ar(Si) / M(SiO2)] × 100%

%Si = (28.086 / 60.084) × 100% = 46.74%

%silicon = 46.74%

%O = 100% - %Si = 100% - 46.74% = 53.26%

%oxygen = 53.26%


Answer: Quartz (SiO2) has a percent composition of 53.26 percent oxygen and 46.74 percent silicon.